OpenStudy (anonymous):

What are the vertical asymptotes of R(x) = 3x - 3 ______ ? x^2 - 4 A) x = 0 B) x = 1, and x = -1 C) x = 2, and x = -2 D) x = 3, and x = -3

OpenStudy (campbell_st):

what values of x make the denominator zero, they are vertical asymptotes so you need to factor the quadratic denominator and then solve for x.

OpenStudy (campbell_st):

You should recognise the denominator is the difference of 2 squares.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the vertical asymptotes are located at points \(x\), such that the denominator is zero, namely: \(x^2-4=0\) so we have to solve this equation: \[\huge {x^2} - 4 = 0\]

OpenStudy (michele_laino):

hint: I can rewrite such equation as follows: \[\huge \left( {x - 2} \right)\left( {x + 2} \right) = 0\] now, please apply the product cancellation law, in order to write the solutions of the above equation

OpenStudy (anonymous):

@Anaise

OpenStudy (anonymous):

stil confused @Michele_Laino

OpenStudy (anaise):

long time no see, rip van pelt ;p

OpenStudy (anonymous):

huh @Anaise can u help me

OpenStudy (michele_laino):

please solve this equations: \(x+2=0\), what is \(x\) ?

OpenStudy (michele_laino):

equation*

OpenStudy (anonymous):

-2

OpenStudy (michele_laino):

that's right! Our first asymptote has this equation: \(x=-2\) Now, please solve this equation: \(x-2=0\), what is \(x\) ?

OpenStudy (anonymous):

so the answer is C

OpenStudy (michele_laino):

that's right!